one pipe fills a tank in 50 minutes another pipe in 75 how long does it take for both pipes together to fill a tanks?

Answer by edjones(8007)

One pipe does 1/50 of the job in 1 minute.
the other pipe doe 1/75 in 1 minute
1/50 + 1/75
=3/150 + 2/150 = 5/150 = 1/30 of the job in 1 minute.
So the tank will be filled in 30 minutes using both pipes.
.
Ed